# A Complete and Bi-Continuous Invariant of Protein Backbones under Rigid Motion

### v2

@@ Abstract
Proteins are large biomolecules that regulate all living organisms and consist of one or several chains. The primary structure of a protein chain is a sequence of amino acid residues whose three main atoms (alpha-carbon, nitrogen, and carbonyl carbon) form a protein backbone. The tertiary structure is the rigid shape of a protein chain represented by atomic positions in 3-dimensional space. Because different geometric structures often have distinct functional properties, it is important to continuously quantify differences in rigid shapes of protein backbones. Unfortunately, many widely used similarities of proteins fail axioms of a distance metric and discontinuously change under tiny perturbations of atoms.
@@ 1 Motivations and the problem statement
A protein is a large biomolecule consisting of one or several chains of amino acid residues. The primary structure (sequence) of a protein chain is a string of residue labels (represented by one or three letters), each denoting one of (usually) 20 standard amino acids [45]. The secondary structure consists of frequent semi-rigid subchains such as $\alpha$-helices and $\beta$-strands [39]. A sequence of a protein is easy to experimentally determine but important functional properties such as interactions with drug molecules depend on a 3-dimensional geometric shape (a tertiary structure or fold) represented by an embedding of all its atoms in $\mathbb{R}^{3}$ [53], see Fig. 1 (left).
@@
Figure 1: Left: all main atoms $N_{i}$, $A_{i}$, $C_{i}$ of a protein chain form a backbone embedded in $\mathbb{R}^{3}$. Middle: each triangle $\triangle N_{i}A_{i}C_{i}$ defines an orthonormal basis ${\bf\it u}_{i},{\bf\it v}_{i},{\bf\it w}_{i}$. The coordinates of the bonds $\overrightarrow{C_{i}N_{i+1}}$, $\overrightarrow{N_{i+1}A_{i+1}}$, $\overrightarrow{N_{i+1}A_{i+1}}$ in this basis form the complete Backbone Rigid Invariant $\mathrm{BRI}$. Right: All rigidly equivalent backbones form a single rigid class. All rigid classes of backbones form the Backbone Rigid Space.
@@
In 1973, Nobel laureate Anfinsen conjectured that the sequence of any protein chain determines its 3D geometric shape [1]. Following this conjecture, neural networks such as AlphaFold2 and RosettaFold [26, 8, 42, 55] optimize millions of parameters to predict a protein fold from its sequence but need re-training [25] on the growing number of experimental structures in the Protein Data Bank (PDB) [11]. The reported accuracies of prediction are based on the LDDT (Local Distance Difference Test) [40, p. 2728] and TM-score [64], which fail the metric axioms. Then clustering algorithms can produce pre-determined clusters and may not be trustworthy [49].
@@
Backbones of the same length (number of residues) can be optimally aligned to minimize the Root Mean Square Deviation (RMSD) between corresponding atoms [19]. This RMSD is slow to compute for all pairs of proteins and gives only distances without mapping the protein universe.
@@
We develop a different approach by mapping the space of protein backbones in analytically defined coordinates similar to geographic-style maps of a new planet. The first question that we should ask about any real data such as protein tertiary structures is “same or different” [51].
@@
Any embedded protein in $\mathbb{R}^{3}$ can be rigidly moved (translated or rotated), which changes all atomic coordinates but the underlying structure remains the same in the sense that different images of a protein under rigid motion have the same properties in a fixed environment. Though proteins are flexible molecules, it is important to distinguish their rigid shapes that can differently interact [20] with other molecules including medical drugs.
@@
Backbones were studied by incomplete invariants such as torsion angles, which allow false positive pairs of non-equivalent backbones $S\not\cong Q$ with $I(S)=I(Q)$. Because all atoms in a backbone $S$ are ordered, their distance matrix determines $S\subset\mathbb{R}^{3}$ up to isometry (any distance-preserving transformation), but is large with a quadratic size in the number $m$ of residues and fails to distinguish mirror images. Adding a sign of orientation creates discontinuity for polygonal chains that are almost mirror-symmetric.
@@
The completeness in 1.2(a) means that $I$ is the strongest possible invariant and hence distinguishes all protein backbones that cannot be exactly matched by rigid motion. The reconstruction in 1.2(b) is more practical because $I$ may not allow an efficiently computable inverse map $I^{-1}$ from an invariant value $I(S)$ to a backbone $S\subset\mathbb{R}^{3}$. The metric axioms for a distance $d$ in 1.2(c) are essential because if the triangle axiom fails with any positive error, results of clustering can be made arbitrary [49].
@@ 2 Past work on similarities of proteins
In the more general context of crystal structures, a canonical description in a reduced unit cell [46] can be achieved by the program TYPIX [47] for inorganic compounds and ACHESYM [30] for macromolecular crystals. Such conventional settings can be considered a complete invariant in the sense of condition (1.2a). However, a reduced cell discontinuously changes under almost any perturbation of atoms, which has been known at least since 1965 [37, p. 80] and was resolved only for generic crystals [58].
@@
The widely used TM-score [64] $\mathrm{TM}=\max\left\{\frac{1}{L_{N}}\sum\limits_{i=1}^{L_{T}}\frac{1}{1+(d_{i}/d_{0})^{2}}\right\}\in[0,1]$ is maximized over all spatial alignments of two backbones, where $\frac{d_{i}}{d_{0}}$ is a normalized distance between aligned atoms, $L_{T}$ is the length (number of atoms) of a template backbone, $L_{N}$ is the length of a given backbone. Since any identical backbones (with all equal coordinates $x,y,z$) have TM-score 1, one way to convert this similarity into a distance is to set $\mathrm{TMD}=1-\mathrm{TM}$ so that $\mathrm{TMD}(S,S)=0$ for any structure $S$. This and many other conversions such as $d=-\log(\mathrm{TM})$ fail the triangle axiom of a metric in 1.2(c). Indeed, let $L_{T}=L_{N}=1$ and $d_{i}/d_{0}$ be pairwise distances $\frac{1}{2},\frac{1}{3},\frac{1}{4}$ between 3 atoms, which satisfy the triangle axiom ($\frac{1}{3}+\frac{1}{4}\geq\frac{1}{2}$). Then $\mathrm{TMD}=1-\mathrm{TM}$ takes the values $\frac{1}{5},\frac{1}{10},\frac{1}{17}$ and fails this axiom ($\frac{1}{10}+\frac{1}{17}<\frac{1}{5}$), and so does $d=-\log(\mathrm{TM})$ with (approximate) values $0.22,0.11,0.06$.
If the triangle axiom fails with any additive error, results of the clustering algorithms $k$-means and DBSCAN can be arbitrarily pre-determined [49]. The authors of another similarity LDDT (Local Distance Difference Test) concluded in [40, p. 2728] that “One disadvantage of the LDDT score is that it does not fulfill the mathematical criteria to be a metric. However, the same is true for most scores”. One metric satisfying all axioms is the Root Mean Square Deviation (RMSD) between optimally aligned ordered atoms [12]. This RMSD is slow to compute for all-vs-all comparisons in the PDB. As a result, many pairs with $\mathrm{RMSD}=0$ remained unnoticed.
If the order of atoms is ignored, the optimal alignment is NP-complete (provably too slow) [36]. Applying random rotations [15] creates many more structures that look different but should be considered rigidly equivalent. This ‘data augmentation’ makes the classification even harder.
The PDB recently implemented a structural superposition [17] of protein backbones by computing the score equal to the sum of absolute values in the upper triangle of the distance-difference matrix (DDM) for the distance matrices between all $\alpha$-carbon atoms $C_{\alpha}$. The description in [17] adds that “to account for possible gaps in the DDMs, caused by a lack of residue coordinates, these scores are multiplied by a scalar between 0-1, where 1 represents the absence of any gaps … low scores represent chains with high structural similarity.” This scaling by values less than 1 likely affects the triangle axiom, which needs checking in the light of the recent reviews [54, 25, 43] of protein folding prediction [26, 38, 42].
@@
Protein backbones are traditionally represented by torsion (dihedral) angles $\varphi_{i},\psi_{i}$ visualized in Ramachandran plots [48]. For a general polygonal line on points $S\subset\mathbb{R}^{3}$, the sequence $\{\phi_{i},\psi_{i}\}$ is invariant under rigid motion but incomplete. Indeed, for any successive points $p_{i},p_{i+1}\in S$, we can shift all points $p_{i+1},\dots,p_{m}$ by a vector $t(\vec{p}_{i+1}-\vec{p}_{i})$ for any $t\neq 0$, which changes the overall rigid shape of $S$ but keeps all relative angles between any straight segments and planes through successive points.
@@
For protein backbones, even if all bond lengths and angles are fixed at ideal values, all torsion angles still should be ordered according to given residues to completely determine the rigid class of a backbone. Even if we keep all torsion angles in order, three invariants per residue cannot uniquely determine a rigid backbone having 3 atoms with 9 coordinates per residue in $\mathbb{R}^{3}$. AlphaFold2 [26] used 6 parameters per residue to define a rigid transformation on every $i$-th triplet (residue triangle) on the main atoms $N_{i},A_{i},C_{i}$ to the next $(i+1)$-st residue triangle. However, the analysis in section 3 will show that rigid shapes of residue triangles substantially vary across the PDB. Our paper strengthens the past approach by defining 9 invariants per each of $m$ residues, which gives $69m-$ 6 invariants in total after subtracting 6 parameters of a global rigid motion in $\mathbb{R}^{3}$.
@@
If we consider a backbone $S$ of $3m$ ordered atoms modulo isometry including reflections, the easier complete invariant known since 1935 [52] is the $3m\times 3m$ matrix $D(S)$ of all pairwise distances whose entry $D_{ij}(S)$ is the Euclidean distance between the $i$-th and $j$-th points of $S$. Any backbone $S$ can be reconstructed from $D(S)$ or, equivalently, from the Gram matrix of scalar products as in [14, Theorem 1], uniquely up to isometry in $\mathbb{R}^{3}$. The matrix $D(S)$ satisfies almost all conditions of Problem 1.2 apart from the linear time/size requirement, which is essential for large proteins.
@@
If a protein is considered a cloud of unordered atoms (ignoring the order along a backbone), such clouds of different sizes can be visualized by eigenvalue invariants (or moments of inertia) characterizing the elongation of the cloud along its principal directions. In 1996, probably the first map of all 4K entries in the PDB appeared in [22, Fig. 5] based on the two largest eigenvalues, see the recent updates in [62, Fig. 2] and PDB-Explorer [24].
@@
In 2020, Holm called for faster visualization of the protein space [23]: “It would be nice to restore the ability to move a lens across fold space in real-time … this ability was based on pre-computed all-against-all structural similarities, which is not manageable with current data”.
@@
In 1977, Kendall [28] started to study configuration spaces of ordered points modulo rigid motion in $\mathbb{R}^{n}$ under the name of size-and-shape spaces [29]. If we consider sequences equivalent also under uniform scaling, the smaller shape space $\Sigma_{2}^{m}$ of $m$ ordered points in $\mathbb{R}^{2}$ can be described as a complex projective space $\mathbb{C}P^{m-1}$ due to the group $\mathrm{SO}(2)$ being identified with the unit circle in the complex space $\mathbb{C}^{1}=\mathbb{R}^{2}$. However, there is no easy description of the space $\Sigma_{3}^{m}$ of $m$-point sequences in $\mathbb{R}^{3}$, which has no multiplicative group structure similar to $\mathbb{R}^{2}=\mathbb{C}^{1}$. This algebraic obstacle prevented a simple solution to Problem 1.2 in dimension $n=3$.
@@ 3 The backbone rigid invariant ($\mathrm{BRI}$)
The $i$-th row of $\mathrm{TRIN}(S)$ uniquely determines the shape of $\triangle N_{i}A_{i}C_{i}$. Many past approaches including AlphaFold2 [26] assumed that all these residue triangles are rigidly equivalent. To test this assumption on the PDB, we filter out unsuitable chains as follows. On May 4, 2024, the PDB had 213,191 entries with 1,091,420 chains. Protocol 3.2 below produced $104,688\approx 49\%$ entries with $707410\approx 65\%$ chains in 4 hours 48 min 11 sec. All experiments were run on CPU Core i7-11700 @2.50GHz RAM 32Gb.
@@
(2) 178153 disordered chains whose some atoms have occupancies $<1$.
@@
Refer to captionRefer to caption
@@
For a backbone of $m$ residues, the first row of the $m\times 9$ matrix $\mathrm{BRI}(S)$ contains only three non-zero coordinates. Hence the matrix $\mathrm{BRI}(S)$ can be considered a vector of length $69(m-1)+3=9m-$ 6. The simplest metric on backbone rigid invariants as vectors in $\mathbb{R}^{9m-6}$ is $L_{\infty}$ equal to the maximum absolute difference between all corresponding coordinates.
@@
(c) Since the complete invariant $\mathrm{BRI}(S)$ of a backbone $S$ is locally defined by determining any $i$-th residue triangle $\triangle N_{i}A_{i}C_{i}$ in the basis of the previous $(i-1)$-st triangle, all rows of the matrix $\mathrm{BRI}(Q)$ for any subchain $Q$ in the full backbone $S$ coincide with the corresponding rows of $\mathrm{BRI}(S)$.

### v3

@@ Abstract
Proteins are large biomolecules that regulate all living organisms and consist of one or several chains. The primary structure of a protein chain is a sequence of amino acid residues whose three main atoms (alpha-carbon, nitrogen, and carbonyl carbon) form a protein backbone. The tertiary structure is the rigid shape of a protein chain represented by atomic positions in 3-dimensional space. Since different geometric structures often have distinct functional properties, it is important to continuously quantify differences in rigid shapes of protein backbones. Unfortunately, many widely used similarities of proteins fail axioms of a distance metric and discontinuously change under tiny perturbations of atoms.
@@ 1 Motivations and the problem statement
A protein is a large biomolecule consisting of one or several chains of amino acid residues. The primary structure (sequence) of a protein chain is a string of residue labels (represented by one or three letters), each denoting one of (usually) 20 standard amino acids [needleman2012protein]. The secondary structure consists of frequent semi-rigid subchains such as $\alpha$-helices and $\beta$-strands [linderstrom1952lane]. A sequence of a protein is easy to experimentally determine but important functional properties such as interactions with drug molecules depend on a 3-dimensional geometric shape (a tertiary structure or fold) represented by an embedding of all its atoms in $\mathbb{R}^{3}$ [scott2017mathematical], see Fig. 1 (left).
@@
Figure 1: Left: all main atoms $N_{i}$, $A_{i}$, $C_{i}$ of a protein chain form a backbone embedded in $\mathbb{R}^{3}$. Middle: each triangle $\triangle N_{i}A_{i}C_{i}$ defines an orthonormal basis ${\bf\it u}_{i},{\bf\it v}_{i},{\bf\it w}_{i}$. The coordinates of the bonds $\overrightarrow{C_{i}N_{i+1}}$, $\overrightarrow{N_{i+1}A_{i+1}}$, $\overrightarrow{A_{i+1}C_{i+1}}$ in this basis form the complete Backbone Rigid Invariant $\mathrm{BRI}$. Right: All rigidly equivalent backbones form a single rigid class. All rigid classes of backbones form the Backbone Rigid Space.
@@
In 1973, Nobel laureate Anfinsen conjectured that the sequence of any protein chain determines its 3D geometric shape [anfinsen1973principles]. Following this conjecture, neural networks such as AlphaFold2 and RosettaFold [jumper2021highly, baek2021accurate, mirdita2022colabfold, van2024fast] optimize millions of parameters to predict a protein fold from its sequence but need re-training [jones2022impact] on the growing number of experimental structures in the Protein Data Bank (PDB) [burley2017protein]. The reported accuracies of prediction are based on the LDDT (Local Distance Difference Test) [mariani2013lddt, p. 2728] and TM-score [zhang2004scoring], which fail the metric axioms. Then clustering algorithms can produce pre-determined clusters and may not be trustworthy [rass2024metricizing].
@@
Backbones of the same length (number of residues) can be optimally aligned to minimize the Root Mean Square Deviation (RMSD) between corresponding atoms [holm2024dali]. This RMSD is slow to compute for all pairs of proteins and gives only distances without mapping the protein universe.
@@
We develop a different approach by mapping the space of protein backbones in analytically defined coordinates similar to geographic-style maps of a new planet. The first question that we should ask about any real data such as protein tertiary structures is “same or different” [sacchi2020same].
@@
Any embedded protein in $\mathbb{R}^{3}$ can be rigidly moved (translated or rotated), which changes all atomic coordinates but the underlying structure remains the same in the sense that different images of a protein under rigid motion have the same properties in a fixed environment. Though proteins are flexible molecules, it is important to distinguish their rigid shapes that can differently interact [heifetz2003effect] with other molecules including medical drugs.
@@
Backbones were studied by incomplete invariants such as torsion angles, which allow false positive pairs of non-equivalent backbones $S\not\cong Q$ with $I(S)=I(Q)$. Since all atoms in a backbone $S$ are ordered, their distance matrix determines $S\subset\mathbb{R}^{3}$ up to isometry (any distance-preserving transformation), but is large with a quadratic size in the number $m$ of residues and fails to distinguish mirror images. Adding a sign of orientation creates discontinuity for polygonal chains that are almost mirror-symmetric.
@@
The completeness in 1.2(a) means that $I$ is the strongest possible invariant and hence distinguishes all protein backbones that cannot be exactly matched by rigid motion. The reconstruction in 1.2(b) is more practical because $I$ may not allow an efficiently computable inverse map $I^{-1}$ from an invariant value $I(S)$ to a backbone $S\subset\mathbb{R}^{3}$. The metric axioms for a distance $d$ in 1.2(c) are essential because if the triangle axiom fails with any positive error, results of clustering can be made arbitrary [rass2024metricizing].
@@ 2 Past work on similarities of proteins
In the more general context of crystal structures, a canonical description in a reduced unit cell [niggli1928krystallographische] can be achieved by the program TYPIX [parthe2013typix] for inorganic compounds and ACHESYM [kowiel2014achesym] for macromolecular crystals. Such conventional settings can be considered a complete invariant in the sense of condition (1.2a). However, a reduced cell discontinuously changes under almost any perturbation of atoms, which has been known at least since 1965 [lawton1965reduced, p. 80] and was resolved only for generic crystals [widdowson2022resolving].
@@
The widely used TM-score [zhang2004scoring] $\mathrm{TM}=\max\left\{\frac{1}{L_{N}}\sum\limits_{i=1}^{L_{T}}\frac{1}{1+(d_{i}/d_{0})^{2}}\right\}\in[0,1]$ is maximized over all spatial alignments of two backbones, where $\frac{d_{i}}{d_{0}}$ is a normalized distance between aligned atoms, $L_{T}$ is the length (number of atoms) of a template backbone, $L_{N}$ is the length of a given backbone. Since any identical backbones (with all equal coordinates $x,y,z$) have TM-score 1, one way to convert this similarity into a distance is to set $\mathrm{TMD}=1-\mathrm{TM}$ so that $\mathrm{TMD}(S,S)=0$ for any structure $S$. This and many other conversions such as $d=-\log(\mathrm{TM})$ fail the triangle axiom of a metric in 1.2(c). Indeed, let $L_{T}=L_{N}=1$ and $d_{i}/d_{0}$ be pairwise distances $\frac{1}{2},\frac{1}{3},\frac{1}{4}$ between 3 atoms, which satisfy the triangle axiom ($\frac{1}{3}+\frac{1}{4}\geq\frac{1}{2}$). Then $\mathrm{TMD}=1-\mathrm{TM}$ takes the values $\frac{1}{5},\frac{1}{10},\frac{1}{17}$ and fails this axiom ($\frac{1}{10}+\frac{1}{17}<\frac{1}{5}$), and so does $d=-\log(\mathrm{TM})$ with (approximate) values $0.22,0.11,0.06$.
If the triangle axiom fails with any additive error, results of the clustering algorithms $k$-means and DBSCAN can be arbitrarily pre-determined [rass2024metricizing]. The authors of another similarity LDDT (Local Distance Difference Test) concluded in [mariani2013lddt, p. 2728] that “One disadvantage of the LDDT score is that it does not fulfill the mathematical criteria to be a metric. However, the same is true for most scores”. One metric satisfying all axioms is the Root Mean Square Deviation (RMSD) between optimally aligned ordered atoms [carugo2003root]. This RMSD is slow to compute for all-vs-all comparisons in the PDB. As a result, many pairs with $\mathrm{RMSD}=0$ remained unnoticed.
If the order of atoms is ignored, the optimal alignment is NP-complete (provably too slow) [lathrop1994protein]. Applying random rotations [draizen2024deep] creates many more structures that look different but should be considered rigidly equivalent. This ‘data augmentation’ makes the classification even harder.
The PDB recently implemented a structural superposition [structural] of protein backbones by computing the score equal to the sum of absolute values in the upper triangle of the distance-difference matrix (DDM) for the distance matrices between all $\alpha$-carbon atoms $C_{\alpha}$. The description in [structural] adds that “to account for possible gaps in the DDMs, caused by a lack of residue coordinates, these scores are multiplied by a scalar between 0-1, where 1 represents the absence of any gaps … low scores represent chains with high structural similarity.” This scaling by values less than 1 likely affects the triangle axiom, which needs checking in the light of the recent reviews [terwilliger2024alphafold, jones2022impact, moore2022protein] of protein folding prediction [jumper2021highly, leman2020macromolecular, mirdita2022colabfold].
@@
Protein backbones are traditionally represented by torsion (dihedral) angles $\varphi_{i},\psi_{i}$ visualized in Ramachandran plots [ramachandran1968conformation]. For a general polygonal line on points $S\subset\mathbb{R}^{3}$, the sequence $\{\phi_{i},\psi_{i}\}$ is invariant under rigid motion but incomplete. Indeed, for any successive points $p_{i},p_{i+1}\in S$, we can shift all points $p_{i+1},\dots,p_{m}$ by a vector $t(\vec{p}_{i+1}-\vec{p}_{i})$ for any $t\neq 0$, which changes the overall rigid shape of $S$ but keeps all relative angles between any straight segments and planes through successive points.
@@
For protein backbones, even if all bond lengths and angles are fixed at ideal values, all torsion angles still should be ordered according to given residues to completely determine the rigid class of a backbone. Even if we keep all torsion angles in order, three invariants per residue cannot uniquely determine a rigid backbone having 3 atoms with 9 coordinates per residue in $\mathbb{R}^{3}$. AlphaFold2 [jumper2021highly] used 6 parameters per residue to define a rigid transformation on every $i$-th triplet (residue triangle) on the main atoms $N_{i},A_{i},C_{i}$ to the next $(i+1)$-st residue triangle. However, the analysis in section 3 will show that rigid shapes of residue triangles substantially vary across the PDB. Our paper strengthens the past approach by defining 9 invariants per each of $m$ residues, which gives 9m69m-6 invariants in total after subtracting 6 parameters of a global rigid motion in $\mathbb{R}^{3}$.
@@
If we consider a backbone $S$ of $3m$ ordered atoms modulo isometry including reflections, the easier complete invariant known since 1935 [schoenberg1935remarks] is the $3m\times 3m$ matrix $D(S)$ of all pairwise distances whose entry $D_{ij}(S)$ is the Euclidean distance between the $i$-th and $j$-th points of $S$. Any backbone $S$ can be reconstructed from $D(S)$ or, equivalently, from the Gram matrix of scalar products as in [dekster1987edge, Theorem 1], uniquely up to isometry in $\mathbb{R}^{3}$. The matrix $D(S)$ satisfies almost all conditions of Problem 1.2 apart from the linear time/size requirement, which is essential for large proteins.
@@
If a protein is considered a cloud of unordered atoms (ignoring the order along a backbone), such clouds of different sizes can be visualized by eigenvalue invariants (or moments of inertia) characterizing the elongation of the cloud along its principal directions. In 1996, probably the first map of all 4K entries in the PDB appeared in [holm1996mapping, Fig. 5] based on the two largest eigenvalues, see the recent updates in [wirth2013protein, Fig. 2] and PDB-Explorer [jin2015pdb].
@@
In 2020, Holm called for faster visualization of the protein space [holm2020dali]: “It would be nice to restore the ability to move a lens across fold space in real-time … this ability was based on pre-computed all-against-all structural similarities, which is not manageable with current data”.
@@
In 1977, Kendall [kendall1977diffusion] started to study configuration spaces of ordered points modulo rigid motion in $\mathbb{R}^{n}$ under the name of size-and-shape spaces [kendall2009shape]. If we consider sequences equivalent also under uniform scaling, the smaller shape space $\Sigma_{2}^{m}$ of $m$ ordered points in $\mathbb{R}^{2}$ can be described as a complex projective space $\mathbb{C}P^{m-1}$ due to the group $\mathrm{SO}(2)$ being identified with the unit circle in the complex space $\mathbb{C}^{1}=\mathbb{R}^{2}$. However, there is no easy description of the space $\Sigma_{3}^{m}$ of $m$-point sequences in $\mathbb{R}^{3}$, which has no multiplicative group structure similar to $\mathbb{R}^{2}=\mathbb{C}^{1}$. This algebraic obstacle prevented a simple solution to Problem 1.2 in dimension $n=3$.
@@ 3 The backbone rigid invariant ($\mathrm{BRI}$)
The $i$-th row of $\mathrm{TRIN}(S)$ uniquely determines the shape of $\triangle N_{i}A_{i}C_{i}$. Many past approaches including AlphaFold2 [jumper2021highly] assumed that all these residue triangles are rigidly equivalent. To test this assumption on the PDB, we filter out unsuitable chains as follows. On May 4, 2024, the PDB had 213,191 entries with 1,091,420 chains. Protocol 3.2 below produced $104,688\approx 49\%$ entries with $707410\approx 65\%$ chains in 4 hours 48 min 11 sec. All experiments were run on CPU Core i7-11700 @2.50GHz RAM 32Gb.
@@
(2) 178153 disordered chains, where some atoms have occupancies <1<1.
@@
Refer to captionRefer to captionRefer to caption
@@
For a backbone of $m$ residues, the first row of the $m\times 9$ matrix $\mathrm{BRI}(S)$ contains only three non-zero coordinates. Hence the matrix $\mathrm{BRI}(S)$ can be considered a vector of length 9(m1)+3=9m69(m-1)+3=9m-6. The simplest metric on backbone rigid invariants as vectors in $\mathbb{R}^{9m-6}$ is $L_{\infty}$ equal to the maximum absolute difference between all corresponding coordinates.
@@
(c) Since the complete invariant $\mathrm{BRI}(S)$ of a backbone $S$ is locally defined by determining any $i$-th residue triangle $\triangle N_{i}A_{i}C_{i}$ in the basis of the previous (i1)(i-1)-st triangle, all rows of the matrix $\mathrm{BRI}(Q)$ for any subchain $Q$ in the full backbone $S$ coincide with the corresponding rows of $\mathrm{BRI}(S)$.